OpenStudy (anonymous):

write as an equivalent radical expression (3x)^-1/3

OpenStudy (anonymous):

\[\sqrt[3]{3x}\]

OpenStudy (anonymous):

Sorry, slight mistake...

jimthompson5910 (jim_thompson5910):

\[\large (3x)^{-\frac{1}{3}}=\frac{1}{(3x)^{\frac{1}{3}}}=\frac{1}{\sqrt[3]{3x}}\] So \[\large (3x)^{-\frac{1}{3}}=\frac{1}{\sqrt[3]{3x}}\]

OpenStudy (anonymous):

Yes, that was the mistake. Forgot to carry out the inverse operation before inserting the radical